OpenStudy (anonymous):

A building is 434 feet tall. For a class project, Jim has to make a scale model of the building to place in the front hall display case. The display case is 19 inches tall. Which scale will allow him to make the tallest model that will fit in the display case?

OpenStudy (anonymous):

a 1 inch = 20 feet b 1 inch = 25 feet c 1 inch = 30 feet

OpenStudy (mertsj):

Let's see, if 1 inch = 20 feet, how much is 2 inches?

OpenStudy (anonymous):

40

OpenStudy (mertsj):

What about 3 inches?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

is the answer a

OpenStudy (mertsj):

Your answer is NOT a

OpenStudy (mertsj):

If 1 inch is 20 feet then 19 inches is 380 feet and the building is taller than 380

OpenStudy (phi):

I would use ratios and test each case: \[ \frac{1 "}{20 \text{ ft}}= \frac{x "}{434 \text{ ft}} \] if you "solve" for x you get the height of the model when 1 " represents 20 feet x= 434/20= 21.7 inches Will this fit ?

OpenStudy (anonymous):

yeah ik I got it wrong

OpenStudy (phi):

try with 1" represents 25 feet: \[ \frac{1 "}{25 \text{ ft}}= \frac{x "}{434 \text{ ft}} \] or \[ x= \frac{434}{25} \] will x be less than 19 inches ?